OpenStudy (anonymous):

how do you write an equation in slope intercept form for the line that satisfies each set of conditions. Find a line that passes through (3,1) perpendicular to the graph of y=-3x+2 did i do it right i got y=1/2x-.5

OpenStudy (mokeira):

@soccerplayer46

OpenStudy (anonymous):

then whats b

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

b what?

OpenStudy (anonymous):

for the equation to be in slope intercept form it has to be y=mx +b

OpenStudy (anonymous):

wouldnt it be that you have to change the slope of \[y=-3x+2\] into a slope of 1/3 and then just plug in (3,1) into the equation \[y=1/3x+b\] it gives you 1=1/3(3)+b so 1=3+b and b=-2 so the final answer is \[y=1/3x-2\]

OpenStudy (anonymous):

@soccerplayer46 does that work for you or do you need a better explanation

OpenStudy (mokeira):

@tiguebaxter is my answer ok though?

OpenStudy (mokeira):

i dont get how you have changed the slope

OpenStudy (anonymous):

the problem with your answer is you didnt account for the perpendicular aspect so you have to find the opposite reciprocal of the original slope given by the first line \[y=-3x+2\] the slope of that line is -3 so the opposite reciprocal is 1/3

OpenStudy (mokeira):

ooo...yeah, i get your point. I have seen my mistake

OpenStudy (anonymous):

alright good im glad i could help @Mokeira and @soccerplayer46